Derive an expression for the effective resistance when three resistors are connected in (i) series (ii) parallel

Grade:Upto college level

1 Answers

Apoorva Arora IIT Roorkee
9 years ago
  1. Series
when three resisters are in series = the current passing through each is the same and the potential differs
let the three resistances beR_{1},R_{2}andR_{3}.
so
V_{1}=IR_{1}
V_{2}=IR_{2}
V_{3}=IR_{3}
the total pot isV=V_{1}+V_{2}+V_{3}
IR=IR_{1}+IR_{2}+IR_{3}
so
R=R_{1}+R_{2}+R_{3}
  1. Parallel
when the arrangement is parallel, the potential difference across each resistance remains the same only the current varies
so
I_{1}=V/R_{1}
I_{2}=V/R_{2}
I_{3}=V/R_{3}
also
I=I_{1}+I_{2}+I_{3}
V/R_{eq}=V/R_{1}+V/R_{2}+V/R_{3}
So
1/R_{eq}=1/R_{1}+1/R_{2}+1/R_{3}
Thanks and Regards
Apoorva Arora
IIT Roorkee
askIITians Faculty